What mass of potassium nitrate is needed to generate 157.0 L of gas, composed of 105.0 L of N2 and 52.0 L of O2 at 0.920 atm and 293 K, using these two reactions?
1
Expert's answer
2020-10-26T14:47:10-0400

First, figure out how moles there are in 105 L and 52 L:


"PV=nRT,\\\space\\
n_{N_2}=\frac{PV}{RT}=4.018\text{ mol},\\\space\\
n_{O_2}=\frac{PV}{RT}=1.990\text{ mol}."


Write the reactions and substitute the products from nitrite to make it one 'combined' reaction useful for calculations:


"2\text{KNO}_3\rightarrow2\text{KNO}_2+\text{O}_2,\\
4\text{KNO}_2\rightarrow2\text{K}_2\text{O}+3\text{O}_2+2\text{N}_2."


Algebraical transformations give us


"4\text{KNO}_3\rightarrow2\text{K}_2\text{O}+5\text{O}_2+2\text{N}_2."


As we see from the reaction, we need two times more nitrate than nitrogen to produce 105 L:


"n_{nitrate/N_2}=2n_{N_2}=8.036\text{ mol}."


Also, 4 molecules of nitrate produce 5 molecules of oxygen:


"n_{nitrate/O_2}=\frac{4}{5}n_{O_2}=2.488\text{ mol}."

The total amount of quantity of potassium nitrate:


"n_{nitrate}=8.036+2.488=10.52\text{ mol}."

What mass of nitrate is required? Multiply it by its molar mass:


"m_{nitrate}=n_{nitrate}\Mu_{nitrate}=\\=10.52\cdot(39+14+16\cdot3)=1063\text{ g}."